Please help me to understand how to solve this word problem.
200 ice cream cones were sold. Double scoops sold for $1.25 and single scoops sold for $1.00. If the total sales for the day was $221.75, how many of each cone were sold.
Thank you

Answer by mananth(16946)

x= number of double scoops
y= number of single scoops
Total comercials
1.00 x + 1.00 y = 200.00 .............1
Total value
1.25 x + 1.00 y = 221.75 .............2
Eliminate y
multiply (1)by -1.00
Multiply (2) by 1.00
-1.00 x -1.00 y = -200.00
1.25 x + 1.00 y = 221.75
Add the two equations
0.25 x = 21.75
/ 0.25
x = 87.00
plug value of x in (1)
1.00 x + 1.00 y = 200.00
87.00 + y = 200.00
y = 200.00 -87.00
y = 113.00
y = 113.00
x= 87.00 number of double scoops
y= 113.00 number of single scoops
m.ananth@hotmail.ca